OpenStudy (anonymous):

How many faces does a right rectangular prism have if it has 8 vertices and 12 edges? A. 4 faces B. 6 faces C. 8 faces D. 10 faces

OpenStudy (dayakar):

V + F = E +2 [ V = vertices = 8, E= edges =12 , F = faces } find F